### v2

# Structures of Moduli Spaces of Generalized Cantor Sets in $(0,1)^{\mathbb{N}}$

Abstract.
For each $\omega\in(0,1)^{\mathbb{N}}$, we may construct a Cantor set $E(\omega)\subset[0,1]$ called a generalized Cantor set for $\omega$. We study the moduli space of $\omega$ denoted by $\mathcal{M}(\omega)\subset(0,1)^{\mathbb{N}}$. It is the set of $\omega^{\prime}$ so that $E(\omega^{\prime})$ is quasiconformally equivalent to $E(\omega)$. In this paper, we show that the set $\mathcal{M}(\omega)$ is measurable in $(0,1)^{\mathbb{N}}$ and we give a necessary condition for $\omega^{\prime}$ to belong to $\mathcal{M}(\omega)$. By using this condition, we show that there are uncountably many moduli spaces in $(0,1)^{\mathbb{N}}$. We also show that except for at most one moduli space, the volume of the moduli space with respect to the standard product measure of $(0,1)^{\mathbb{N}}$ vanishes.
Key words and phrases:
Cantor set, quasiconformal mapping
2010 Mathematics Subject Classification:
Primary 30C62; Secondary 30F25
1. Introduction and main results
Let $\mathcal{C}$ be a Cantor set in the complex plane $\mathbb{C}$. It is a totally disconnected perfect set in $\mathbb{C}$. We consider the moduli space $\mathcal{M}_{\mathcal{C}}$ of $\mathcal{C}$, the set of Cantor sets $\mathcal{C}^{\prime}$ which are quasiconformally equivalent to $\mathcal{C}$. Here, we say that a Cantor set $\mathcal{C}^{\prime}$ in $\mathbb{C}$ is quasiconformally equivalent to $\mathcal{C}$ if there exists a quasiconformal mapping $\varphi:{\mathbb{C}}\to{\mathbb{C}}$ such that $\varphi(\mathcal{C}^{\prime})=\mathcal{C}$.
In a series of papers ([9], [10], [11]) , we have studied the quasiconformal equivalence of Cantor sets and their moduli spaces. For example, we show that the limit set of a Schottky group and Cantor Julia sets of hyperbolic rational functions belong to the moduli space of the standard middle one third Cantor set $\mathcal{C}_{0}$. Moreover, we show that the limit set of a generalized Schottky group with a rank one parabolic does not belong to the moduli space of $\mathcal{C}_{0}$. Recently, R. Hidalugo ([5]) shows that the number of the moduli spaces of Cantor sets generated by finitely generated Kleinian groups is at most four.
In this paper, we consider generalized Cantor sets $E(\omega)$ which are defined by $\omega=(q_{n})_{n=1}^{\infty}\in\Omega:=(0,1)^{\mathbb{N}}$. The construction of $E(\omega)$ is a generalization of that of $\mathcal{C}_{0}$ as follows. To construct $E(\omega)$ for $\omega=(q_{n})_{n=1}^{\infty}\in(0,1)^{\mathbb{N}}$, we start with the closed unit interval $I=[0,1]$. At the first step, we remove an open interval $J_{1}^{1}$ of length $q_{1}(=q_{1}|I|)$ from $I$ so that $I\setminus J_{1}^{1}$ consists of two closed intervals $I_{1}^{1}$ and $I_{1}^{2}$ of the same length. We put $E_{1}(\omega)=I_{1}^{1}\cup I_{1}^{2}$. We repeat this operation inductively. Namely, for $E_{n-1}(\omega)=\cup_{j=1}^{2^{n-1}}I_{n-1}^{j}$, we remove open intervals $J_{n}^{2j-1}$ of the length $q_{n}|I_{n-1}^{j}|$ from $I_{n-1}^{j}$ so that $I_{n-1}^{j}\setminus J_{n}^{2j-1}$ consists of two closed intervals $I_{n}^{2j-1}$ and $I_{n}^{2j}$ of the same length $(j=1,\dots,2^{n-1})$. Thus, we obtain $E_{n}(\omega)=\cup_{j=1}^{2^{n}}I_{n}^{j}$ and $I\setminus E_{n}(\omega)=\cup_{j=1}^{2^{n}-1}J_{n}^{j}$, where $J_{n}^{2j}=J_{n-1}^{j}$. Then, $E(\omega):=\cap_{n=1}^{\infty}E_{n}(\omega)$ is a Cantor set and we call it the generalized Cantor set for $\omega=(q_{n})_{n=1}^{\infty}\in\Omega$. Therefore, the set of generalized Cantor sets is identified with $\Omega$.
For a given $\omega\in\Omega$, we consider the moduli space of $\omega$ denoted by $\mathcal{M}(\omega)$. It is the set of $\omega^{\prime}\in\Omega$ such that $E(\omega^{\prime})$ is quasiconformally equivalent to $E(\omega)$. In this paper, we study the structure of $\mathcal{M}(\omega)$ in $\Omega$. First, we show the following:
Theorem I.
Under the standard product topology of $\Omega$ $(=(0,1)^{\mathbb{N}})$, the moduli space $\mathcal{M}(\omega)$ is an $F_{\sigma}$ set. In particular, it is a Borel measurable set in $\Omega$.
The space $\Omega$ has the standard probability measure $m_{\infty}$ defined by the Lebesgue measure on $(0,1)$. Since each moduli space $\mathcal{M}(\omega)$ for $\omega\in\Omega$ is measurable, we may consider the volume $m_{\infty}(\mathcal{M}(\omega))$ of $\mathcal{M}(\omega)$.
Theorem II.
Except for at most one moduli space, the volume of any moduli space with respect to $m_{\infty}$ is zero.
We consider conditions for $\omega^{\prime}$ to belong to $\mathcal{M}(\omega)$.
Theorem III.
Let $\omega=(q_{n})_{n=1}^{\infty}$ and $\omega^{\prime}=(q_{n}^{\prime})_{n=1}^{\infty}$ be in $\Omega$. Suppose that both $\omega$ and $\omega^{\prime}$ are increasing. Then $\omega^{\prime}\in\mathcal{M}(\omega)$ only if
(1.1) $\sup_{n\in\mathbb{N}}\left|\log\frac{\log(1-q_{n})}{\log(1-q_{n}^{\prime})}\right|<\infty.$
By using this result, we may show the following:
Theorem IV.
There are uncountably many moduli spaces of generalized Cantor sets in $\Omega$.
The result may be interesting if it is compared with the fact by R. Hidalugo mentioned above, namely, the finiteness of the number of Moduli spaces of Cantor sets generated by finitely generated Kleinian groups.
2. Preliminaries
2.1. Infinite sequences in $(0,1)$ and generalized Cantor sets
As we have mentioned in §1, the set of generalized Cantor sets is identified with $\Omega:=(0,1)^{\mathbb{N}}$.
Definition 2.1.
Let $\omega=(q_{n})_{n=1}^{\infty}\in\Omega:=(0,1)^{\mathbb{N}}$. We say that $\omega$ has a lower bound and an upper bound if $q_{n}\geq\delta$ and $q_{n}\leq\delta$ $(n\in\mathbb{N})$ for some $\delta\in(0,1)$ independent of $n$, respectively.
We have already obtained a sufficient condition for two elements with a lower bound to be quasiconformally equivalent to each other.
Theorem 2.1 (cf. [10] Theorem II, Corollary 1.2 and Theorem III).
Let $\omega=(q_{n})_{n=1}^{\infty},\omega^{\prime}=(q_{n}^{\prime})_{n=1}^{\infty}$ be in $\Omega$ with a lower bound. Then $E(\omega^{\prime})$ is quasiconformally equivalent to $E(\omega)$ if $\sup_{n\in\mathbb{N}}\left|\log\frac{1-q_{n}}{1-q_{n}^{\prime}}\right|<\infty$. In particular, if $\omega$ has lower and upper bounds, then the Cantor set $E(\omega)$ is quasiconformally equivalent to the standard middle one third Cantor set $\mathcal{C}_{0}$.
Moreover, if $\omega$ does not have an upper bound, then $E(\omega)$ is not quasiconformally equivalent to $\mathcal{C}_{0}$.
Definition 2.2.
We say that $\omega=(q_{n})_{n=1}^{\infty}\in\Omega$ is increasing if $q_{n}\leq q_{n+1}<1$ for any $n\in\mathbb{N}$.
For $\omega=(q_{n})_{n=1}^{\infty}\in\Omega$, we take $E_{k}(\omega)=\cup_{j=1}^{2^{k}}I_{k}^{j}$ which is the closed set at the $k$-step of the construction given in the introduction. Here, $I_{k}^{j}$ is a closed interval located at the left of $I_{k}^{j-1}$ $(j=1,2,\dots,2^{k}-1)$. The set $0,1]\setminus E_{k}(\omega)$ consists of $2^{k}-1$ open intervals $J_{k}^{1},\dots,J_{k}^{2^{k}-1}$. The interval $J_{k}^{j}$ is located between $I_{k}^{j}$ and $I_{k}^{j+1}$.
Now, we calculate the lengths of $I_{k}^{j}$ and $J_{k}^{j}$ for the argument in §4.
Because of the construction of $E_{k}(\omega)$, we have
$|I_{k}^{j}|=\frac{1}{2}(1-q_{k})|I_{k-1}^{j}|=2^{-k}\prod_{i=1}^{k}(1-q_{i})\quad(j\in\{1,\dots,2^{k}\}).$
Viewing the construction of $E_{k}(\omega)$ from $E_{k-1}(\omega)$, we see that the open interval $I_{k-1}^{j}$ is divided into $I_{k}^{2j-1},J_{k}^{2j-1}$ and $I_{k}^{2j}$. Hence, if $j$ is odd, then
(2.1) $|J_{k}^{j}|=q_{k}|I_{k-1}^{j}|=\frac{2q_{k}}{1-q_{k}}|I_{k}^{j}|.$
When $j$ is even, we put $j=2^{\ell}m$ for some integer $\ell$ with $1\leq\ell\leq k-1$ and some odd number $m$. Since $J_{k}^{j}$ is located between $I_{k}^{j}$ and $I_{k}^{j+1}$, we see that $J_{k}^{j}=J_{k-1}^{j/2}=J_{k-1}^{2^{\ell-1}}$. Repeating this argument, we have $J_{k}^{j}=J_{k-\ell}^{m}$. Since $m$ is odd, we have
(2.2) $|J_{k}^{j}|$ $=$ $|J_{k-\ell}^{m}|=\frac{2q_{k-\ell}}{1-q_{k-\ell}}|I_{k-\ell}^{m}|=\frac{2q_{k-\ell}}{1-q_{k-\ell}}\cdot 2^{-k+\ell}\prod_{i=1}^{k-\ell}(1-q_{i})$
$=$ $\frac{2^{\ell+1}q_{k-\ell}}{\prod_{i=k-\ell}^{k}(1-q_{i})}|I_{k}^{j}|.$
2.2. Topology of $\Omega$ and sequences of the complements of Cantor sets.
As we have mentioned in §2, the set of generalized Cantor sets is identified with $\Omega:=(0,1)^{\mathbb{N}}$. We consider the space $\Omega$ with the weak topology (cf. [3]). That is, a sequence $\omega_{m}=(q_{n}^{m})_{n=1}^{\infty}$ $(m=1,2,\dots)$ converges to $\omega=(q_{n})_{n=1}^{\infty}$ if and only if $\lim_{m\to\infty}q_{n}^{m}=q_{n}$ for each $n\in\mathbb{N}$.
For a sequence $\{\omega_{m}\}$ converging to $\omega$, we consider the complements $X(\omega_{m})$ of $E(\omega_{m})$ $(m\in\mathbb{N})$ and the Carathéodory kernel $N(\{X(\omega_{m})\})_{m=1}^{\infty}$ of the sequence of the domains. Here, $z\in\widehat{\mathbb{C}}$ belongs to the kernel $N(\{X(\omega_{m}\}_{m=1}^{\infty})$ if and only if there exist a neighborhood $U$ of $z$ and $M\in\mathbb{N}$ such that $U\subset X(\omega_{m})$ for all $m>M$.
Proposition 2.1.
Suppose that $\{\omega_{m}\}_{m=1}^{\infty}\subset\Omega$ converges to $\omega$ in $\Omega$ as $m\to\infty$. Then, $N(\{X(\omega_{m})\}_{m=1}^{\infty})=X(\omega)$.
Proof.
If $z\not\in\mathbb{R}$ or $z=\infty$, then obviously, $z\in N(\{X(\omega_{m})\}_{m=1}^{\infty}$. Also, it is obvious that $z\in N(\{X(\omega_{m})\}_{m=1}^{\infty})$ if $z\in(-\infty,0)\cup(1,+\infty)$.
Let $z$ be in $0,1]\setminus E(\omega)$. Then, there exists $k\in\mathbb{N}$ such that $z\in[0,1]\setminus E_{k}(\omega)$. Since $E_{k}(\omega)=\cup_{j=1}^{2^{k}}I_{k}^{j}$, the point $z$ belongs to $I_{k}^{J}$ for some $J\in\{1,2,\dots,2^{k}\}$. Take a sufficiently small neighborhood $U$ of $z$ so that $U\cap\mathbb{R}\subset I_{k}^{J}$. Let $I_{k,m}^{J}$ be the interval of $E_{k}(\omega_{m})$ corresponding to $I_{k}^{J}$. The intervals $I_{k}^{J}$ and $I_{k,m}^{J}$ depend only on $(q_{n})_{n=1}^{k}$ and $(q_{n}^{m})_{n=1}^{k}$, and $q_{n}^{m}\to q_{n}$ for $n=1,\dots k$. Hence, $I_{k,m}^{J}$ is close to $I_{k}^{J}$ if $m$ is sufficiently large. Therefore, we see that $U\cap\mathbb{R}\subset I_{k,m}^{J}\subset X(\omega_{m})$ for a sufficiently large $m$ and we are convinced that $X(\omega)\subset N(\{X(\omega_{m})\}_{m=1}^{\infty})$.
For $x\in E(\omega)$, take a neighborhood $U$ of $z$. From the construction of $E(\omega)$, there exist $k\in\mathbb{N}$, $J\in\{1,2,\dots,2^{k}\}$ and an interval $I_{k}^{J}$ of $E_{k}(\omega)$ such that $U\supset I_{k}^{J}$. Because of the same reason as above, we may find an interval $I_{k,m}^{J}$ for a large $M\in\mathbb{N}$ such that $U\supset I_{k,M}^{J}$. This implies that $U\not\subset X(\omega_{M})$.
For any $m>M$, there exists an interval $I_{k,m}^{j}$ such that $I_{k,m}^{j}\subset I_{k,M}^{J}$. Thus, we see that $U\supset I_{k,m}^{j}$ and $U\not\subset X(\omega_{m})$ for $m>M$. So, we conclude that $X(\omega)=N(\{X(\omega_{m})\}_{m=1}^{\infty})$ and the proof is completed. ∎
For a sequence $\varphi_{n}:G\to G_{n}$ of $K$-quasiconformal mappings $(n=1,2,\dots)$, the Carathéodory kernel $N(\{G_{n}\}_{n=1}^{\infty})$ becomes the image of the limit mapping if the limit mapping is also a quasiconformal mapping.
Theorem 2.2 ([7] Chapter II, Theorem 5.4 ).
Let $G\subset\widehat{\mathbb{C}}$ be a domain with at least two boundary points. Suppose that $K$-quasiconformal mappings $\varphi_{n}:G\to G_{n}$ converges to a quasiconformal mapping $\varphi$ locally uniformly on $G$ as $n\to\infty$. Then $\varphi$ maps $G$ onto a component of $N(\{G_{n}\}_{n=1}^{\infty})$.
2.3. The canonical pants decomposition (cf. [10] §3)
We put $X(\omega)=\widehat{\mathbb{C}}\setminus E(\omega)$ for $\omega=(q_{n})_{n=1}^{\infty}\in\Omega:=(0,1)^{\mathbb{N}}$. The surface $X(\omega)$ is a hyperbolic Riemann surface of infinite type. We define a pants decomposition of $X(\omega)$ associated to $\omega$.
First, we put $\gamma_{0}:=\{ix\mid x\in\mathbb{R}\}\cup\{\infty\}$. It is a simple closed geodesic in $X(\omega)$.
Next, we consider $E_{2}(\omega)=\cup_{j=1}^{4}I_{2}^{j}$ and take circles $C_{2}^{j}$ passing through two open intervals adjacent to $I_{2}^{j}$ $(j=1,\dots,4)$. Then, the geodesics $\gamma_{2}^{j}$ in $X(\omega)$ homotopic to $C_{2}^{j}$ together with $\gamma_{0}$ make two pairs of pants, $P_{1}^{1}$ and $P_{1}^{2}$ so that $\partial P_{1}^{1}=\gamma_{0}\cup\gamma_{2}^{1}\cup\gamma_{2}^{2}$ and $\partial P_{1}^{2}=\gamma_{0}\cup\gamma_{2}^{3}\cup\gamma_{2}^{4}$.
Inductively, we construct a pair of pants $P_{k}^{j}$ for $k\in\mathbb{N}$ and $j\in\{1,2,\dots 2^{k}\}$. Each $P_{k}^{j}$ is bounded by geodesics $\gamma_{k}^{j},\gamma_{k+1}^{2j-1}$ and $\gamma_{k+1}^{2j}$, and $\{P_{k}^{j}\}_{k\in\mathbb{N},j=1,2,\dots,2^{k}}$ gives a pants decomposition of $X(\omega)$. We call it the canonical pants decomposition of $X(\omega)$.
2.4. Extremal length and the modulus of a ring domain.
Let $W$ be a domain in $\widehat{\mathbb{C}}$ (or a Riemann surface) and $\Gamma$ a family of curves $\gamma$ in $W$. We consider a measurable conformal metric $\rho=\rho(z)|dz|$ on $W$ with
$A(\rho):=\iint_{W}\rho^{2}(z)dxdy\in(0,+\infty).$
Then, the extremal length $\lambda(\Gamma)$ of $\Gamma$ is defined by
$\lambda(\Gamma)=\sup_{\rho}\frac{L_{\Gamma}(\rho)^{2}}{A(\rho)},$
where $L_{\Gamma}(\rho)=\inf_{\gamma\in\Gamma}\int_{\gamma}\rho(z)|dz|$ and the supremum is taken over all above conformal metrics $\rho$. Obviously, the extremal length is a conformal invariant.
A domain $W$ in $\widehat{\mathbb{C}}$ is called a ring domain if $\partial W$ consists of two connected components. The modulus of a ring domain $W$ denoted by $\textrm{Mod}(W)$ is defined by
$\textrm{Mod}(W)=\lambda(\Gamma_{W})^{-1},$
where $\Gamma_{W}$ is the set of closed curves in $W$ separating the two connected components of $\partial W$.
A Jordan domain $Q$ together with four points located anti-clockwisely on $\partial Q$, say $x_{1},x_{2},x_{3}$ and $x_{4}$, is called a quadrilateral. We denote the quadrilateral by $(Q;x_{1},x_{2},x_{3},x_{4})$ or simply by $Q$ if it does not cause any confusion. We denote by $\overline{xy}$ for $x,y\in\partial Q$ the sub-arc of $\partial Q$ from $x$ to $y$. Then, the modulus of $(Q;x_{1},x_{2},x_{3},x_{4})$ denoted by $\textrm{Mod}(Q;x_{1},x_{2},x_{3},x_{4})$ or $\textrm{Mod}(Q)$ is defined by
$\textrm{Mod}(Q;x_{1},x_{2},x_{3},x_{4})=\textrm{Mod}(Q)=\lambda(\Gamma_{Q}),$
where $\Gamma_{Q}$ is the set of curves connecting $\overline{x_{1}x_{2}}$ and $\overline{x_{3}x_{4}}$ in $Q$.
It is known that both $\lambda(\Gamma_{W})$ and $\lambda(\Gamma_{Q})$ above have extremal metrics which attain the supremums of the definition of the extremal length (cf. [1]).
[…]
Lemma 2.1.
$\lambda(\Gamma_{A(1,R)})=2\lambda(\Gamma_{A(1,R)\cap\mathbb{H}})$.
Proof.
We have
$\lambda(\Gamma_{A(1,R)})=\frac{L_{\Gamma_{A(1,r)}}(\rho_{0})^{2}}{A(\rho_{0})}$
and
[…]
Since
[…]
we obtain the desired result. ∎
Finally, we exhibit a theorem by Maskit on the comparison of the hyperbolic length and the extremal length.
Theorem 2.3 (Maskit [8]).
Let $\ell_{R}(\gamma)$ be the hyperbolic length of a closed geodesic $\gamma$ in a hyperbolic Riemann surface $R$ and $\lambda(\Gamma)$ the extremal length of the curve family $\Gamma$ of closed curve homotopic to $\gamma$ in $R$. Then,
(2.3) $\lambda(\Gamma)\leq\frac{1}{2}\ell_{R}(\gamma)e^{\ell_{R}(\gamma)/2}$
holds.
3. Proof of Theorem I
For $\omega\in\Omega$, we consider a function $F_{\omega}:\Omega\to\mathbb{R}_{\geq 1}\cup\{+\infty\}$ by
$F_{\omega}(\omega^{\prime})=\begin{cases}\inf_{\varphi}K(\varphi),&\omega^{\prime}\in\mathcal{M}(\omega)\\ +\infty,&\omega^{\prime}\not\in\mathcal{M}(\omega),\end{cases}$
where the infimum is taken over all quasiconformal mappings $\varphi:\mathbb{C}\to\mathbb{C}$ with $\varphi(E(\omega))=E(\omega^{\prime})$ and $K(\varphi)$ is the maximal dilatation of $\varphi$.
Lemma 3.1.
$F_{\omega}$ is a lower semi-continuous function on $\Omega$.
Proof.
Take $\omega_{0}\in\Omega$ and we show that $F_{\omega}$ is lower semi-continuous at $\omega_{0}$. Since $F_{\omega}(\omega_{0})=+\infty$ if $\omega_{0}\not\in\mathcal{M}(\omega)$, we may assume that $\omega_{0}\in\mathcal{M}(\omega)$.
For $\omega_{0}=(q_{n}^{0})_{n=1}^{\infty}\in\mathcal{M}(\omega)$, we consider $\omega_{m}=(q_{n}^{m})_{n=1}^{\infty}$ converging to $\omega_{0}$ as $m\to\infty$.
Since $F_{\omega}(\omega_{0})<+\infty$, we may assume that $K:=\liminf_{m\to\infty}F_{\omega}(\omega_{m})<+\infty$. Therefore, for any $\varepsilon>0$, there exists $N\in\mathbb{N}$ such that $K(\varphi_{m})\leq K+\varepsilon$ for any $m\in\mathbb{N}$ with $m>N$, where $\varphi_{m}:\mathbb{C}\to\mathbb{C}$ $(m=1,2,\dots)$ are extremal quasiconformal mappings with $\varphi(E(\omega_{0}))=E(\omega_{m})$. Hence, $K(\varphi_{m})=F_{\omega}(\omega_{m})$. Therefore, $\{\varphi_{m}\}_{m=1}^{\infty}$ is a normal family.
Let $\varphi$ be a limit map of a subsequence of the family. There are the following three possibilities of $\varphi$ ([7] Chapter II).
  1. (1)
    $\varphi\equiv\infty$;
  2. (2)
    $\varphi\equiv z_{0}$ for some $z_{0}\in\mathbb{C}$;
  3. (3)
    $\varphi$ is a quasiconformal self map of $\mathbb{C}$.
Since the convergence to $\varphi$ is locally uniform on $\mathbb{C}$ and $\varphi_{m}(E(\omega))=E(\omega_{m})$, we see that only the case (3) is possible. Thus, $K(\varphi)\leq K+\varepsilon$ and we obtain
$K(\varphi)\leq K=\liminf_{m\to\infty}F_{\omega}(\omega_{m}).$
On the other hand, since $\omega_{m}\to\omega_{0}$, we have seen in Proposition 2.1 that the Carathéodory kernel $N(\{E(\omega_{m})^{c}\}_{m=1}^{\infty})$ of the sequence domains $\{E(\omega_{m})^{c}\}_{m=1}^{\infty}$ is $E(\omega_{0})^{c}$. It follows from Theorem 2.2 that $\lim_{m\to\infty}\varphi_{m}(E(\omega)^{c})=\varphi(E(\omega)^{c})=E(\omega_{0})^{c}$ and $\varphi(E(\omega))=E(\omega_{0})$. Therefore, we obtain
$F_{\omega}(\omega_{0})\leq K(\varphi)\leq K=\liminf_{m\to\infty}F_{\omega}(\omega_{m})$
and we verify that $F_{\omega}$ is lower semi-continuous at $\omega_{0}$ as desired. ∎
Now, the proof of Theorem I is obvious. Indeed, from the definition of $F_{\omega}$, we have
$\mathcal{M}(\omega)=F_{\omega}^{-1}([1,+\infty))=\cup_{n=1}^{\infty}F_{\omega}^{-1}([1,n]).$
Since $F_{\omega}$ is lower semi-continuous, $F_{\omega}^{-1}([1,n])$ $(n=1,2,\dots)$ are closed sets in $\Omega$. Hence, $\mathcal{M}(\omega)$ is an $F_{\sigma}$ set. This completes the proof of Theorem I.
4. Proof of Theorem II
4.1. The shift map and quasiconformal equivalence
Let $\sigma:\Omega\to\Omega$ be the shift map. Namely, for $\omega=(q_{n})_{n=1}^{\infty}\in\Omega$,
$\sigma(\omega)=(q_{n+1})_{n=1}^{\infty}=(q_{2},q_{3},\dots).$
The shift map $\sigma$ is a classical object in the ergodic theory. Actually, the followings are well-known.
Proposition 4.1.
  1. (1)
    The shift map $\sigma$ is measure preserving, that is, $m_{\infty}(\sigma^{-1}(A))=m_{\infty}(A)$ for any measurable set $A\subset\Omega$.
  2. (2)
    The action of $\sigma:\Omega\to\Omega$ is ergodic. Namely, if a measurable set $A$ in $\Omega$ is invariant under the action of $\sigma$, then $m_{\infty}(A)=0$ or $1$.
Now, we establish a relationship between the shift map and the quasiconformal equivalence.
Lemma 4.1.
Let $\omega_{i}=(q_{n}^{(i)})_{n=1}^{\infty}$ be in $\Omega$ ($i=1,2$). If $\sigma(\omega_{1})$ and $\sigma(\omega_{2})$ are quasiconformally equivalent to each other, then so are $\omega_{1}$ and $\omega_{2}$. In particular, if $\omega,\omega^{\prime}\in\sigma^{-1}(\alpha)$ for some $\alpha\in\Omega$, then $\mathcal{M}(\omega)=\mathcal{M}(\omega^{\prime})$.
Proof.
Suppose that $\sigma(\omega_{2})$ is quasiconformally equivalent to $\sigma(\omega_{1})$. Then, there exists a quasiconformal mapping $\varphi:\mathbb{C}\to\mathbb{C}$ with $\varphi(E(\sigma(\omega_{1})))=E(\sigma(\omega_{2}))$.
We take $W(\sigma(\omega_{i})):=X(\sigma(\omega_{i}))\cap\Delta(2^{-1},\{(1-q_{1}^{(i)})q_{2}^{(i)}\}^{-1}-2^{-1})$ ($i=1,2$), where $X(\omega)=\widehat{\mathbb{C}}\setminus E(\omega)$ for $\omega\in\Omega$ and $\Delta(a,r)$ is an open disk of radius $r>0$ centered at $a\in\mathbb{C}$. From the construction, $W(\sigma(\omega_{i}))$ is conformally equivalent to $W_{i}:=X(\omega_{i})\cap\Delta(4^{-1}(1-q_{1}^{(i)})q_{2}^{(i)},2^{-1}(1-2^{-1}(1-q_{1}^{(i)})q_{2}^{(i)}))$ and $W_{i}^{\prime}:=X(\omega_{i})\cap\Delta(1-4^{-1}(1-q_{1}^{(i)})q_{2}^{(i)},2^{-1}(1-2^{-1}(1-q_{1}^{(i)})q_{2}^{(i)}))$. Let $F_{i}$ and $F_{i}^{\prime}$ be a conformal mapping from $W(\sigma(\omega_{i}))$ to $W_{i}$ and $W_{i}^{\prime}$, respectively $(i=1,2)$.
Since $\varphi$ is a homeomorphism with $\varphi(E(\sigma(\omega_{1})))=E(\sigma(\omega_{2}))$, there exist Jordan domains $U_{1},U_{2}$ such that $E(\sigma(\omega_{1}))\subset U_{1}\subset W(\sigma(\omega_{1}))$, $E(\sigma(\omega_{2}))\subset U_{2}\subset W(\sigma(\omega_{2}))$ and $U_{2}=\varphi(U_{1})$. Then, $\psi_{1}:=F_{2}\circ\varphi\circ{F_{1}|_{F_{1}(U_{1})}}^{-1}:F_{1}(U_{1})\to F_{2}(U_{2})$ and $\psi_{1}^{\prime}:=F_{2}^{\prime}\circ\varphi\circ{F_{1}^{\prime}|_{F_{1}^{\prime}(U_{1})}}^{-1}:F_{1}^{\prime}(U_{1})\to F_{2}^{\prime}(U_{2})$ give a quasiconformal mapping $\Phi$ from $F_{1}(U_{1})\cup F_{1}^{\prime}(U_{1})$ to $F_{2}(U_{2})\cup F_{2}^{\prime}(U_{2})$ with $\Phi(E(\omega_{1}))=E(\omega_{2})$. We have also a quasiconformal mapping $\Psi:\widehat{\mathbb{C}}\setminus\overline{F_{1}(U_{1})\cup F_{1}^{\prime}(U_{1})}\to\widehat{\mathbb{C}}\setminus\overline{F_{2}(U_{2})\cup F_{2}^{\prime}(U_{2})}$ with $\Psi(\partial F_{1}(U_{1}))=\partial F_{2}(U_{2})$ and $\Psi(\partial F_{1}^{\prime}(U_{1}))=\partial F_{2}^{\prime}(U_{2})$ since both regions are annuli.
Here, we use the following fact on the gluing of quasiconformal mappings.
Proposition 4.2.
Let $X,Y$ be Riemann surfaces. We consider simple closed curves $\alpha\subset X$ and $\beta\subset Y$ with $X\setminus{\alpha}=X_{1}\sqcup X_{2}$ and $Y\setminus\beta=Y_{1}\sqcup Y_{2}$, respectively. Suppose that there exist quasiconformal mappings $f_{i}:X_{i}\to Y_{i}$ $(i=1,2)$ such that $f_{1}(\alpha)=f_{2}(\alpha)=\beta$. Then, there exist an annular neighborhood $A$ of $\alpha$ and a quasiconformal mapping $f$ on $A$ into $Y$ such that
  1. (1)
    $A^{\prime}:=f(A)$ is an annular neighborhood of $\beta$;
  2. (2)
    we put
    (4.1) $F(p)=\begin{cases}f_{i}(p),&p\in X_{i}\setminus A\quad(i=1,2)\\ f(p),&p\in U.\end{cases}$
    Then, $F$ is a quasiconformal mapping from $X$ to $Y$.
By applying this Proposition for $f_{1}=\Phi,f_{2}=\Psi$, we verify that $\omega_{1}$ and $\omega_{2}$ are quasiconformally equivalent to each other. ∎
Notation. We define $\mathcal{M}(\sigma^{-1}(\omega))$ by $\mathcal{M}(\omega^{\prime})$ for some $\omega^{\prime}\in\sigma^{-1}(\omega)$. From the above lemma, $\mathcal{M}(\sigma^{-1}(\omega))$ depends on $\omega$ but it is independent of $\omega^{\prime}\in\sigma^{-1}(\omega)$.
Corollary 4.1.
For $\omega=(q_{k})_{k=1}^{\infty}\in\Omega$ and for any $n\in\mathbb{N}$,
(4.2) $\sigma^{n}(\mathcal{M}(\omega))\supset\mathcal{M}(\sigma^{n}(\omega)),\quad\mathcal{M}(\sigma^{-n}(\omega))\supset\sigma^{-n}(\mathcal{M}(\omega)).$
Proof.
[…]
The same argument gives another inclusion of (4.2). Indeed, if $\widehat{\omega}=(\widehat{q}_{k})_{k=1}^{\infty}$ belongs to $\sigma^{-n}(\mathcal{M}(\omega))$, then there exist $\omega^{\prime}=(q_{k}^{\prime})_{k=1}^{\infty}\in\mathcal{M}(\omega)$ and $(x_{1},\dots,x_{n})\in(0,1)^{n}$ such that
$\widehat{\omega}=(x_{1},\dots,x_{n},q_{1}^{\prime},q_{2}^{\prime},\dots).$
Hence, $\sigma^{n}(\widehat{\omega})=\omega^{\prime}$ is quasiconformally equivalent to $\omega$ and we conclude that $\widehat{\omega}\in\mathcal{M}(\sigma^{-n}(\omega))$ from Lemma 4.1. ∎
4.2. Proof of the theorem
It is not hard to see that for a subset $S\subset\Omega$, $\sigma(\sigma^{-1}(S))=S$ and $\sigma^{-1}(\sigma(S))\supset S$, but $\sigma^{-1}(\sigma(S))\not=S$ in general. However, it is valid for moduli spaces.
Lemma 4.2.
For any $n\in\mathbb{N}$,
(4.3) $\sigma^{-n}(\sigma^{n}(\mathcal{M}(\omega)))=\mathcal{M}(\omega).$
Proof.
It suffices to show that $\sigma^{-n}(\sigma^{n}(\mathcal{M}(\omega)))\subset\mathcal{M}(\omega)$.
Let $\widehat{\omega}$ be in $\sigma^{-n}(\sigma^{n}(\mathcal{M}(\omega)))$. Then, there exist $\omega^{\prime}=(q_{k}^{\prime})_{k=1}^{\infty}\in\mathcal{M}(\omega)$ and $(x_{1},\dots,x_{n})\in(0,1)^{n}$ such that
$\widehat{\omega}=(x_{1},\dots,x_{n},q_{n+1}^{\prime},q_{n+2}^{\prime},\dots).$
Hence, we have $\sigma^{n}(\widehat{\omega})=\sigma^{n}(\omega^{\prime})$. Thus, we conclude that $\widehat{\omega}$ is quasiconformally equivalent to $\omega^{\prime}\in\mathcal{M}(\omega)$. So, $\widehat{\omega}$ belongs to $\mathcal{M}(\omega)$ and we have $\sigma^{-n}(\sigma^{n}(\mathcal{M}(\omega)))\subset\mathcal{M}(\omega)$. ∎
Since $\sigma$ is measure preserving,
$m_{\infty}(\sigma^{-n}(\mathcal{M}(\omega)))=m_{\infty}(\mathcal{M}(\omega)),$
for any $n\in\mathbb{N}$. Also, we have from (4.3)
$m_{\infty}(\sigma^{n}(\mathcal{M}(\omega)))=m_{\infty}(\mathcal{M}(\omega)),$
for any $n\in\mathbb{N}$.
Hence, if $m_{\infty}(\mathcal{M}(\omega))>0$, then
$m_{\infty}(\sigma^{-n}(\mathcal{M}(\omega))\cap\mathcal{M}(\omega))>0$
holds for all $n\in\mathbb{N}$ except at most finitely many $n$. Since $\mathcal{M}(\sigma^{-n}(\omega))\supset\sigma^{-n}(\mathcal{M}(\omega))$ from (4.2), we have
$m_{\infty}(\mathcal{M}(\sigma^{-n}(\omega))\cap\mathcal{M}(\omega))>0.$
Thus, we conclude $\mathcal{M}(\sigma^{-n}(\omega))=\mathcal{M}(\omega)$ because moduli spaces are equivalence classes. This means $\mathcal{M}(\omega)=\mathcal{M}(\sigma^{-n}(\omega))\supset\sigma^{-n}(\mathcal{M}(\omega))$. Hence, we see that $\mathcal{M}(\omega)=\sigma^{-n}(\mathcal{M}(\omega))$ as measurable sets. Therefore, $m_{\infty}(\mathcal{M}(\omega))=1$ from the ergodicity of $\sigma$ and we conclude that at most one moduli space has the positive volume, if any.
5. Proof of Theorem III
We may assume that $\sup_{n\in\mathbb{N}}q_{n}=1$ and $\sup_{n\in\mathbb{N}}q_{n}^{\prime}=1$. Indeed, if $\sup_{n\in\mathbb{N}}q_{n}<1$ and $E(\omega^{\prime})$ is quasiconformally equivalent to $E(\omega)$, then it follows from Theorem 2.1 that $\sup_{n\in\mathbb{N}}q_{n}^{\prime}<1$. Indeed, since $\omega$ has both upper and lower bounds, $E(\omega)$ is quasiconformally equivalent to the standard middle one third Cantor set $\mathcal{C}_{0}$. However, if $\omega^{\prime}$ does not have an upper bound, then $E(\omega^{\prime})$ is not quasiconformally equivalent to $\mathcal{C}_{0}$ nor $E(\omega)$. This is a contradiction. Therefore, the condition (2) is satisfied in this case. The same argument also works when $\sup_{n\in\mathbb{N}}q_{n}^{\prime}=1$ So, we assume that $\sup_{n\in\mathbb{N}}q_{n}=1$ and $\sup_{n\in\mathbb{N}}q_{n}^{\prime}=1$.
Suppose that $E(\omega^{\prime})$ is quasiconformally equivalent to $E(\omega)$ but the condition (1.1) is not satisfied, namely
(5.1) $\sup_{n\in\mathbb{N}}\left|\log\frac{\log(1-q_{n})}{\log(1-q_{n}^{\prime})}\right|=\infty.$
Step 1 : Analyzing the canonical pants decomposition (Part I).
Let $\{P_{k}^{j}\}_{j=1,2,\dots,2^{k};k\in\mathbb{N}}$ (resp. $\{{P^{\prime}}_{k}^{j}\}_{j=1,2,\dots,2^{k};k\in\mathbb{N}}$) be the canonical pants decomposition of $X(\omega)$ (resp. $X(\omega^{\prime})$). Each $P_{k}^{j}$ (resp. ${P^{\prime}}_{k}^{i}$) is bounded by geodesics $\gamma_{k}^{j},\gamma_{k+1}^{2j-1}\mbox{ and }\gamma_{k+1}^{2j}$ (resp. $\gamma_{k}^{j},\gamma_{k+1}^{2j-1}\mbox{ and }\gamma_{k+1}^{2j}$). Then, Kinjo ([6]) shows the following:
Lemma 5.1 (Kinjo).
  1. (1)
    If $j=1$ or $2^{k}$, then
    $\ell_{X(\omega)}(\gamma_{k}^{j})\leq\frac{2\pi^{2}}{\log\frac{1+q_{k}}{1-q_{k}}};$
  2. (2)
    If $j=2^{\ell}m$ or $2^{\ell}m+1$ for some $\ell\in\{1,\dots,k-1\}$ and an odd number $m$, then
    $\ell_{X(\omega)}(\gamma_{k}^{j})\leq\max\left\{\frac{2\pi^{2}}{\log\frac{1+q_{k}}{1-q_{k}}},\frac{2\pi^{2}}{\log\frac{1-q_{k}+2q_{k-\ell}}{1-q_{k}}}\right\},$
where $\ell_{X(\omega)}(c)$ is the hyperbolic length of a curve $c$ in $X(\omega)$.
Since both $\omega=(q_{n})_{n=1}^{\infty}$ and $\omega^{\prime}=(q_{n}^{\prime})_{n=1}^{\infty}$ are increasing, we may assume that $q_{n},q_{n}^{\prime}>\delta$ $(n=1,2,\dots)$ for some $\delta\in(0,1)$. Hence, we obtain
$\frac{1+q_{k}}{1-q_{k}},\frac{1-q_{k}+2q_{k-\ell}}{1-q_{k}}\geq 1+\frac{2\delta}{1-q_{k}}$
and
(5.2) $\ell_{X(\omega)}(\gamma_{k}^{j})\leq\frac{2\pi^{2}}{\log\left(1+\frac{2\delta}{1-q_{k}}\right)}\leq\frac{2\pi^{2}}{\log\left(1+\frac{2\delta}{1-\delta}\right)}.$
Similarly, we have
(5.3) $\ell_{X(\omega^{\prime})}({\gamma^{\prime}}_{k}^{j})\leq\frac{2\pi^{2}}{\log\left(1+\frac{2\delta}{1-q_{k}^{\prime}}\right)}\leq\frac{2\pi^{2}}{\log\left(1+\frac{2\delta}{1-\delta}\right)}.$
In particular, they are bounded above and converge to zero as $n\to\infty$ because $\lim_{n\to\infty}q_{n}=\lim_{n\to\infty}q^{\prime}_{n}=1$. From the hyperbolic trigonometry, we see that short geodesics do not pass through any pair of pants in the canonical pants decomposition.
Lemma 5.2.
There exists a constant $d=d(\delta)$ depending only on $\delta$ such that the hyperbolic length of any geodesic curve $c$ in $P_{k}^{j}$ or ${P^{\prime}}_{k}^{j}$ $(k\in\mathbb{N};j\in\{1,2,\dots,2^{k}\})$ connecting two points on the boundary curves is not less than $d$.
Proof.
It is known that a pair of pants is decomposed into congruent right-angled hexagons. Then, we see that an application of the argument of the proof of Theorem 7.19.2 in [2] §7.19 confirms the statement. ∎
Step 2 : Analyzing the canonical pants decomposition (Part II : Extremal length).
We estimate $\ell_{X(\omega)}(\gamma_{k}^{j})$ and $\ell_{X(\omega^{\prime})}(\gamma_{k}^{j})$ from below.
When $j$ is even, $I_{k}^{j}$ is located between $J_{k}^{j-1}$ and $J_{k}^{j}$. As we have seen in (2.1) and (2.2),
$|J_{k}^{j-1}|=\frac{2q_{k}}{1-q_{k}}|I_{k}^{j}|\leq\frac{2}{1-q_{k}}|I_{k}^{j}|$
and
$|J_{k}^{j}|=\frac{2^{\ell+1}q_{k-\ell}}{\prod_{i=k-\ell}^{k}(1-q_{i})}|I_{k}^{j}|\leq 2\left(\frac{2}{1-q_{k}}\right)^{k}|I_{k}^{j}|,$
since $(q_{n})_{n=1}^{\infty}$ is increasing. We put $J_{k}^{j-1}=(x_{1},x_{2}),I_{k}^{j}=[x_{2},x_{3}]$ and $J_{k}^{j}=(x_{3},x_{4})$ $(x_{1}<x_{2}<x_{3}<x_{4})$.
We consider a four-punctured sphere $X_{k,j}:=\widehat{\mathbb{C}}\setminus\cup_{i=1}^{4}x_{i}$ and we denote by $\Gamma_{0}$ the set of closed curves on $X_{k,j}$ which are homotopic to $\gamma_{k}^{j}$ in $X_{k,j}$. For each $\gamma_{0}\in\Gamma_{0}$, $\bar{\gamma}$ is the image of $\gamma$ by the reflection with respect to the real line and $\overline{\Gamma}_{0}=\{\bar{\gamma}\mid\gamma\in\Gamma_{0}\}$. We also put $\gamma^{+}=\{\gamma\cup(\overline{\gamma\cap\mathbb{L}})\}\setminus(\gamma\cap\mathbb{L})$ and $\Gamma_{0}^{+}=\{\gamma^{+}\mid\gamma\in\Gamma_{0}\}$, where $\mathbb{L}$ is the lower half plane in $\mathbb{C}$. Since $\Gamma_{0}=\overline{\Gamma_{0}}$, it follows from a symmetric principle that
$\lambda(\Gamma_{0})=\frac{1}{2}\lambda(\Gamma_{0}^{+}),$
where $\lambda(\Gamma)$ stands for the extremal length of a curve family $\Gamma$.
We denote by $\widetilde{\Gamma}_{0}$ the set of curves in the upper half plane $\mathbb{H}$ connecting two intervals $(x_{1},x_{2})$ and $(x_{3},x_{4})$. For each $\gamma\in\Gamma_{0}^{+}$, there exists $\tilde{\gamma}\in\widetilde{\Gamma}_{0}$ such that $\tilde{\gamma}\subset\gamma$. Therefore, we see
$\lambda(\widetilde{\Gamma}_{0})\leq\lambda(\Gamma_{0}^{+})$
from the definition of the extremal length.
Furthermore, from the symmetric property, we have
(5.4) $\lambda(\widehat{\Gamma}_{0})=2\lambda(\widetilde{\Gamma}_{0}),$
where $\widehat{\Gamma}_{0}$ is the set of closed curves separating $(-\infty,x_{1}]\cup[x_{4},+\infty)\cup\{\infty\}$ from $x_{2},x_{3}]$. Indeed, it is known that there exists a conformal mapping $f:\mathbb{H}\to A(1,R)\cap\mathbb{H}$ for some $R>1$ such that $f(x_{1})=-R,f(x_{2})=-1,f(x_{3})=1$ and $f(x_{4})=R$, where $A(1,R)=\{z\in\mathbb{C}\mid 1<|z|<R\}$. Since both $\mathbb{H}$ and $A(1,R)\cap\mathbb{H}$ are Jordan domains, it follows from Carathéodory’s theorem that $f$ is extended to the boundaries as a homeomorphism. Then, we define a mapping $F$ from a ring domain $A_{k,j}:=\widehat{\mathbb{C}}\setminus((-\infty,x_{1}]\cup[x_{4},+\infty)\cup\{\infty\}\cup[x_{2},x_{3}])$ onto $A(1,R)$ by
$F(z)=\begin{cases}f(z),\quad z\in\mathbb{H}\cup(-x_{1},x_{2})\cup(x_{3},x_{4})\\ \overline{f(\bar{z})},\quad z\in\mathbb{L}.\end{cases}$
From the reflection principle, the mapping $F$ is a conformal mapping from $A_{k,j}$ onto $A(1,R)$ with $F((\mathbb{H},x_{1},x_{2},x_{3},x_{4}))=(A(1,R)\cap\mathbb{H},-R,-1,1,R))$. Since the extremal length is a conformal invariant, we have
$\lambda(\widehat{\Gamma}_{0})=\lambda(\Gamma_{A(1,R)}),\quad\lambda(\widetilde{\Gamma}_{0})=\lambda(\Gamma_{A(1,R)\cap\mathbb{H}})$
and we obtain (5.4) from Lemma 2.1.
Thus, we conclude
(5.5) $\lambda(\Gamma_{0})\geq\frac{1}{4}\lambda(\widehat{\Gamma}_{0}).$
On the other hand,
(5.6) $\lambda(\widehat{\Gamma}_{0})=(\textrm{Mod}(A_{k,j}))^{-1}.$
We also see that the ring domain $A_{k,j}$ is conformally equivalent to a ring domain $B_{k,j}:=\widehat{\mathbb{C}}\setminus([-\infty,-\alpha]\cup[0,1]\cup[\beta,+\infty])$, where $\alpha=|J_{k}^{j-1}|/|I_{k}^{j}|\leq 2/(1-q_{k})$ and $\beta=1+|J_{k}^{j}|/|I_{k}.j|\leq 1+2(2/(1-q_{k}))^{k}$. Therefore,
$\textrm{Mod}(A_{k,j})=\textrm{Mod}(B_{k,j}).$
For $\alpha^{\prime}=2/(1-q_{k})$ and $\beta^{\prime}=1+2(2/(1-q_{k}))^{k}=1+2\alpha^{\prime k}$, the ring domain $A_{k,j}^{\prime}:=\widehat{\mathbb{C}}\setminus([-\infty,-\alpha^{\prime}]\cup[0,1]\cup[\beta^{\prime},+\infty])$ contains $B_{j,k}$ and we obtain
$\textrm{Mod}(B_{k,j})\leq\textrm{Mod}(A_{k,j}^{\prime}).$
Furthermore, by taking a Möbius transformation, we see that the domain is conformally equivalent to a Teichmüller domain $A_{j,k}^{\prime\prime}:=\mathbb{C}\setminus([-1,0]\cup[\alpha^{\prime}(\beta^{\prime}-1)/(\alpha^{\prime}+\beta^{\prime}),+\infty)$ and
$\textrm{Mod}(A_{k,j}^{\prime})=\textrm{Mod}(A_{k,j}^{\prime\prime}).$
Now, we introduce a function $\Psi(t)$ which represents the modulus of the Teichmüller domain $T(t):=\mathbb{C}\setminus([-1,0]\cup[t,+\infty)$ for $t>0$. Actually, $\textrm{Mod}(T(t))$, the modulus of $T(t)$ is $\log\Psi(t)$. As for $\Psi(t)$, the following is known:
Proposition 5.1 ([4]).
For $t>0$, $\Psi(t)$ is a strictly increasing function and
$t+1\leq(\sqrt{t+1}+\sqrt{t})^{2}\leq\Psi(t)\leq 4(\sqrt{t+1}+\sqrt{t})^{2}\leq 16(t+1).$
Therefore, we have
$\textrm{Mod}(A_{k,j}^{\prime\prime})=\textrm{Mod}(T(\alpha^{\prime}(\beta^{\prime}-1)/(\alpha^{\prime}+\beta^{\prime})))\leq\log\left\{16\left(\frac{\alpha^{\prime}(\beta^{\prime}-1)}{\alpha^{\prime}+\beta^{\prime}}+1\right)\right\}.$
In fact, since $\beta^{\prime}=2\alpha^{\prime k}+1$, we have
$\frac{\alpha^{\prime}(\beta^{\prime}-1)}{\alpha^{\prime}+\beta^{\prime}}+1$ $=$ $\frac{2\alpha^{\prime k+1}+2\alpha^{\prime k}+\alpha^{\prime}+1}{2\alpha^{\prime k}+\alpha^{\prime}+1}$
$=$ $\alpha^{\prime}+1-\frac{\alpha^{\prime 2}+\alpha}{2\alpha^{\prime k}+\alpha^{\prime}+1}\leq\alpha^{\prime}+1.$
and
$\log\left\{16\left(\frac{\alpha^{\prime}(\beta^{\prime}-1)}{\alpha^{\prime}+\beta^{\prime}}+1\right)\right\}\leq 4\log 2+\log(\alpha^{\prime}+1).$
Hence, we have
$\textrm{Mod}(A_{k,j}^{\prime\prime})\leq 4\log 2+\log\left(\frac{2}{1-q_{k}}+1\right)\leq 6\log 2-\log(1-q_{k}).$
Thus, we obtain
(5.7) $\textrm{Mod}(A_{k,j})\leq\textrm{Mod}(A_{k,j}^{\prime\prime})\leq 6\log 2-\log(1-q_{k})$
and
$\lambda(\Gamma_{0})\geq\frac{1}{4}\lambda(\widehat{\Gamma}_{0})=\left(4\textrm{Mod}(A_{k,j})\right)^{-1}\geq\frac{\pi/2}{6\log 2-\log(1-q_{k})}.$
It follows from (2.3) that
$2\lambda(\Gamma_{0})\leq\ell_{X_{k,j}}(\gamma_{0})\exp\{\ell_{X_{k,j}}(\gamma_{0})/2\},$
where $\gamma_{0}$ is the hyperbolic geodesic in $\Gamma_{0}$. Since $X_{k,j}\supset X(\omega)$, we obtain
$\ell_{X_{k,j}}(\gamma_{0})\leq\ell_{X(\omega)}(\gamma_{k}^{j})$
and
$2\lambda(\Gamma_{0})\leq\ell_{X_{k,j}}(\gamma_{0})\exp\{\ell_{X_{k,j}}(\gamma_{0})/2\}\leq\ell_{X(\omega)}(\gamma_{k}^{j})\exp\{\ell_{X(\omega)}(\gamma_{k}^{j})/2\}.$
On the other hand, we have seen that
$\frac{1}{2}\ell_{X(\omega)}(\gamma_{k}^{j})\leq\frac{\pi^{2}}{\log\left(1+\frac{2\delta}{1-\delta}\right)}=:C(\delta).$
Therefore, we have
$\ell_{X(\omega)}(\gamma_{k}^{j})\geq 2e^{-C(\delta)}\lambda(\Gamma_{0})\geq\frac{e^{-C(\delta)}\pi}{6\log 2-\log(1-q_{k})}.$
Similarly, we have
$\ell_{X(\omega^{\prime})}({\gamma^{\prime}}_{k}^{j})\geq\frac{e^{-C(\delta)}\pi}{6\log 2-\log(1-q^{\prime}_{k})}..$
When $j$ is odd, we see that the above argument still works and we obtain the same estimates as above.
Thus, we get the following estimates.
(5.8) $\frac{e^{-C(\delta)}\pi}{6\log 2-\log(1-q_{k})}\leq\ell_{X(\omega)}(\gamma_{k}^{j})\leq\frac{2\pi^{2}}{\log\left(1+\frac{2\delta}{1-q_{k}}\right)},$
and
(5.9) $\frac{e^{-C(\delta)}\pi}{6\log 2-\log(1-q^{\prime}_{k})}\leq\ell_{X(\omega)}({\gamma^{\prime}}_{k}^{j})\leq\frac{2\pi^{2}}{\log\left(1+\frac{2\delta}{1-q_{k}^{\prime}}\right)}$
for each $j\in\{1,2,\dots,2^{k}\}$.
We compare the left-hand side and the right-hand side of each other.
$R(q_{k},q_{k}^{\prime})$ $:=$ $\frac{2\pi e^{C(\delta)}(6\log 2-\log(1-q^{\prime}_{k}))}{\log\left(1+\frac{2\delta}{1-q_{k}}\right)}$
$=$ $2\pi e^{C(\delta)}\frac{\log(1-q_{k}^{\prime})}{\log(1-q_{k})}\cdot\frac{\frac{6\log 2}{\log(1-q_{k}^{\prime})}-1}{\frac{\log(2\delta+1-q_{k})}{\log(1-q_{k})}-1}.$
and
$R(q_{k}^{\prime},q_{k})$ $:=$ $\frac{2\pi e^{C(\delta)}(6\log 2-\log(1-q_{k}))}{\log\left(1+\frac{2\delta}{1-q_{k}^{\prime}}\right)}$
$=$ $2\pi e^{C(\delta)}\frac{\log(1-q_{k})}{\log(1-q_{k}^{\prime})}\cdot\frac{\frac{6\log 2}{\log(1-q_{k}^{\prime})}-1}{\frac{\log(2\delta+1-q_{k}^{\prime})}{\log(1-q_{k}^{\prime})}-1}.$
Thus, we have the following :
Lemma 5.3.
If $q_{k},q_{k}^{\prime}\geq\delta>0$, then $\inf_{k\in\mathbb{N}}R(q_{k},q_{k}^{\prime})>0$ if and only if
$\inf_{k\in\mathbb{N}}\frac{\log(1-q_{k}^{\prime})}{\log(1-q_{k})}>0$
and $\inf_{k\in\mathbb{N}}R(q_{k}^{\prime},q_{k})>0$ if and only if
$\inf_{k\in\mathbb{N}}\frac{\log(1-q_{k})}{\log(1-q_{k}^{\prime})}>0.$
Remark 5.1.
E. Kinjo ([6]) has estimated $\ell_{X(\omega)}(\gamma_{k}^{j})$ from below. Actually, she shows
$\ell_{X(\omega)}(\gamma_{k}^{j})\geq 2\sinh^{-1}\left(1/\sinh\left(\frac{\pi^{2}}{\log\left(\frac{1+q_{k}}{2q_{k}}\right)}\right)\right).$
She has used a different method from ours, while the above estimate is weaker than ours.
Step 3 : Looking at the images of geodesics by quasiconformal maps.
Since we are supposing (5.1), there exists a subsequence $\{n_{k}\}_{k=1}^{\infty}$ such that $\lim_{k\to\infty}q_{n_{k}}=\lim_{k\to\infty}q^{\prime}_{n_{k}}=1$. In particular, $\lim_{k\to\infty}\ell({\gamma^{\prime}}_{n_{k}}^{j})=0$ by (5.9).
Suppose that there exists a $K$-quasiconformal mapping $\varphi:\mathbb{C}\to\mathbb{C}$ with $\varphi(E(\omega^{\prime}))=E(\omega)$. By Wolpert’s lemma ([12]), we have
$\ell_{X(\omega)}([\varphi({\gamma}^{\prime}_{n_{k}})])\leq K\ell_{X(\omega^{\prime})}({\gamma^{\prime}}_{n_{k}}^{j}),$
where $\varphi(\gamma_{n_{k}}^{j})]$ is the geodesic in $X(\omega)$ homotopic to $\varphi({\gamma^{\prime}}_{n_{k}}^{j})$.
Hence, $\lim_{k\to\infty}\ell_{X(\omega)}([\varphi([\gamma_{n_{k}}^{j}])=0$ and $\ell_{X(\omega)}([\varphi([\gamma_{n_{k}}^{j}])<d$ if $k$ is sufficiently large, where $d>0$ is a constant given in Lemma 4.2. Therefore, the geodesic $\varphi({\gamma^{\prime}}_{n_{k}}^{j})]$ cannot pass through any pair of pants of the pants decomposition of $X(\omega)$. Thus, we obtain:
Lemma 5.4.
If $k$ is sufficiently large, then $\varphi({\gamma^{\prime}}_{n_{k}}^{j})]$ is a boundary curve of a pair of pants of the canonical pants decomposition of $X(\omega)$.
Step 3 : Comparing the lengths of geodesics.
From (5.1), the following two cases occurs.
$\mbox{Case I : }\quad\frac{\log(1-q_{n_{k}})}{\log(1-q^{\prime}_{n_{k}})}\to 0\quad(k\to\infty)$
or
$\mbox{Case II : }\quad\frac{\log(1-q^{\prime}_{n_{k}})}{\log(1-q_{n_{k}})}\to 0\quad(k\to\infty).$
Case I. From Lemma 5.3, $R(q_{n_{k}}^{\prime},q_{n_{k}})<K^{-1}$ for a sufficiently large $k$. This implies that
$\ell_{X(\omega)}([\varphi({\gamma^{\prime}}_{n_{k}}^{j})])$ $\leq$ $K\ell_{X(\omega^{\prime})}({\gamma^{\prime}}_{n_{k}}^{j})\leq\frac{2\pi^{2}K}{\log\left(1+\frac{2\delta}{1-q_{n_{k}}^{\prime}}\right)}$
$=$ $KR(q^{\prime}_{n_{k}},q_{n_{k}})\cdot\frac{e^{C(\delta)}\pi}{6\log 2-\log(1-q_{k})}$
$\leq$ $KR(q^{\prime}_{n_{k}},q_{n_{k}})\ell_{X(\omega)}(\gamma_{n_{k}}^{m})<\ell_{X(\omega)}(\gamma_{n_{k}}^{m})$
for any $j,m\in\{1,2,\dots,2^{n_{k}}\}$.
Moreover, since $(q_{n})_{n\in\mathbb{N}}$ is increasing, for any $n<n_{k}$, $q_{n}\leq q_{n_{k}}$ and
$0<-\log\left(1-q_{n}\right)\leq-\log\left(1-q_{n_{k}}\right)$
holds. Therefore, we also obtain
$\ell_{X(\omega)}([\varphi({\gamma^{\prime}}_{n_{k}}^{j})])$ $\leq$ $KR(q^{\prime}_{n_{k}},q_{n_{k}})\cdot\frac{e^{C(\delta)}\pi}{6\log 2-\log\left(1-q_{n_{k}}\right)}$
$\leq$ $KR(q^{\prime}_{n_{k}},q_{n_{k}})\cdot\frac{e^{C(\delta)}\pi}{6\log 2+\log\left(1-q_{n}\right)}<\ell_{X(\omega)}(\gamma_{n}^{m})$
for $n<n_{k}$ and $m\in\{1,2,\dots,2^{n}\}$.
Therefore, we see that the geodesic $\varphi({\gamma^{\prime}}_{n_{k}}^{j})]$, which is a boundary of a pair of pants in the canonical pants decomposition for $\omega$, has to be in the inside of a topological disk bounded by $\gamma_{n_{k}}^{i}$ of some $i\in\{1,2,\dots,2^{n_{k}}\}$.
Case II. From Lemma 5.3, $R(q_{n_{k}},q_{n_{k}}^{\prime})<K^{-1}$ for a sufficiently large $k$. By using the same argument as in Case I above, we obtain
$\ell_{X(\omega)}([\varphi({\gamma^{\prime}}_{n_{k}}^{j})])>\ell_{X(\omega)}(\gamma_{n_{k}}^{m})$
for any $j,m\in\{1,2,\dots,2^{n_{k}}\}$, and
$\ell_{X(\omega)}([\varphi({\gamma^{\prime}}_{n_{k}}^{j})])>\ell_{X(\omega)}(\gamma_{n_{k}}^{m})$
for $n>n_{k}$ and $m\in\{1,2,\dots,2^{n}\}$.
Therefore, we see that the geodesic $\varphi({\gamma^{\prime}}_{n_{k}}^{j})]$, has to be in the outside of any topological disk bounded by $\gamma_{n_{k}}^{i}$ $(i\in\{1,2,\dots,2^{n_{k}}\})$.
Step 4 : Using a topological argument to get a contradiction. Let $D_{n_{k},j}$ (resp.  $D^{\prime}_{n_{k},j}$) be the disk bounded by $\gamma_{n_{k}}^{j}$ (resp. ${\gamma^{\prime}}_{n_{k}}^{j}$). We put $W_{n_{k},j}:=\widehat{\mathbb{C}}\setminus\cup_{j=1}^{2^{n_{k}}}\overline{D_{n_{k},j}}$ and $W^{\prime}_{n_{k},j}:=\widehat{\mathbb{C}}\setminus\cup_{j=1}^{2^{n_{k}}}\overline{D^{\prime}_{n_{k},j}}$. Both $W_{n_{k},j}$ and $W^{\prime}_{n_{k},j}$ are planar domains bounded by mutually disjoint $2^{n_{k}}$ simple closed curves and
$W_{n_{k},j}\cap E(\omega)=\emptyset,\quad W^{\prime}_{n_{k},j}\cap E(\omega^{\prime})=\emptyset.$
Since $\varphi:\mathbb{C}\to\mathbb{C}$ is a homeomorphism and $\varphi(E(\omega^{\prime}))=E(\omega)$, the image $\varphi(W^{\prime}_{n_{k},j})$ is also a domain bounded by $2^{n_{k}}$ simple closed curves $\varphi({\gamma^{\prime}}_{n_{k}}^{j})$ $(j=1,2,\dots,2^{n_{k}})$ and $\varphi(W^{\prime}_{n_{k},j})\cap E(\omega)=\emptyset$.
Let $\widehat{W}_{n_{k},j}$ be a domain bounded by $\varphi({\gamma^{\prime}}_{n_{k}}^{j})]$ $(j=1,2,\dots,2^{n_{k}})$. Since $\varphi({\gamma^{\prime}}_{n_{k}}^{j})]$ is homotopic to $\varphi({\gamma^{\prime}}_{n_{k}}^{j})$, $\widehat{W}_{n_{k},j}$ has the same property as that of $\varphi(W^{\prime}_{n_{k},j})$. Namely, $\widehat{W}_{n_{k},j}\cap E(\omega)=\emptyset$. Moreover, $\varphi({\gamma^{\prime}}_{n_{k}}^{j})]$ is a boundary of a pair of pants of the canonical pants decomposition of $X(\omega)$ for a sufficiently large $k$ (Lemma 5.4). Now, let us consider Case I in Step 3.
As we have seen there, each geodesic $\varphi({\gamma^{\prime}}_{k}^{j})]$ is in the inside of some $D_{n_{k},i}$ $(1\leq i,j\leq 2^{n_{k}})$. Furthermore, $\varphi({\gamma^{\prime}}_{k}^{j})]\not=\partial D_{n_{k},i}=\gamma_{n_{k}}^{i}$ because $\ell_{X(\omega)}([\varphi({\gamma^{\prime}}_{n_{k}}^{j})])<\ell_{X(\omega)}(\gamma_{n_{k}}^{i})$. Hence, the simply connected domain $D_{n_{k},i}$ contains at least two geodesics in $\partial\widehat{W}_{n_{k},j}$, otherwise $\widehat{W}_{n_{k},j}\cap E(\omega)\not$$\emptyset$. This is absurd since both $W_{n_{k},j}$ and $\widehat{W}_{n_{k},j}$ are bounded by the same number of simple closed curves and they do not include any point in $E(\omega)$.
We may use the same argument for Case II and have a contradiction as well. Thus, we have completed the proof of Theorem III.
6. Proof of Theorem IV
The proof of Theorem IV is done by giving an example.
For each $\alpha\in(1,+\infty)$, we put
$q_{n}(\alpha)=1-\exp\left(-{n}^{\alpha}\right)$
and $\omega(\alpha)=(q_{n}(\alpha))_{n=1}^{\infty}\in\Omega$. Then, $\omega(\alpha)$ is increasing and $\lim_{n\to\infty}q_{n}(\alpha)=1$. Furthermore, for $1\leq\alpha^{\prime}<\alpha<+\infty$,
$\frac{\log(1-q_{n}(\alpha))}{\log(1-q_{n}(\alpha^{\prime}))}=\left({n}\right)^{\alpha-\alpha^{\prime}}\to\infty\quad(n\to\infty).$
Hence, $\mathcal{M}(\omega(\alpha))\not=\mathcal{M}(\alpha^{\prime})$ from Theorem III and we conclude that $\{\mathcal{M}(\alpha)\}_{\alpha\in(1,+\infty)}$ is uncountable.